OpenStudy (anonymous):

-6q + 5.8 = 23.8 I really need help asap

pooja195 (pooja195):

Subtract 5.8 from both sides 23.8-5.8= ?

OpenStudy (anonymous):

17. somethin

OpenStudy (anonymous):

im sorry im rewally bad at math

OpenStudy (anonymous):

oh its 18 flat

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

oh yay pooja is back its 18

pooja195 (pooja195):

Good :-) −6q=18 Now simply divide both sides by -6 -6/-6=1 so just p p=18/-6 18/-6=?

OpenStudy (anonymous):

what does the / mean?

pooja195 (pooja195):

Divide

OpenStudy (anonymous):

3 sorry

OpenStudy (anonymous):

-3

pooja195 (pooja195):

haha it's ok :-) Yes correct so q=-3

OpenStudy (anonymous):

-3 is one of the answers thank you so much

pooja195 (pooja195):

yw ^_^